OpenStudy (anonymous):

Find all the zeros of the equation. -4x^4 - 44x^2 + 3600 = 0 A. 5, -5, 6i, -6i B. 5, -5, 6i, 0 C. -5, -6i

OpenStudy (phi):

you could say x^2 = y and replace x^4 with (x^2)^2 = y^2 now solve -4y^2 -44y+3600=0 I would simplify by dividing all terms by -4 (on both sides of the = sign) y^2+11y-900=0 now use the quadratic formula to find y or factor (knowing that 25*36= 900) you get 2 y values. but x= ±sqrt(y), so you get 2 x's for each y, or 4 answers